OpenStudy (anonymous):

divide and simplify your answers to lowest terms 1/3 divided by 5/7

OpenStudy (anonymous):

\[\frac{\frac{1}{3}}{\frac{5}{7}} = \frac{1}{3} * \frac{7}{5} = \frac{7}{15} \]

OpenStudy (anonymous):

thank you...